A company produces steel rods. The lengths of the steel rods are normally distributed with a mean of 136.9-cm and a standard deviation of 2.2-cm. For shipment, 29 steel rods are bundled together.

Find the probability that the average length of a randomly selected bundle of steel rods is between 137.8-cm and 138.1-cm.
P(137.8-cm < M < 138.1-cm) =

Enter your answer as a number accurate to 4 decimal places. Answers obtained using exact z-scores or z-scores rounded to 3 decimal places are accepted.
Answer #1

We are taking a sample of 29 rods out of the population with

Population Mean(\mu) 136.9, Population Standard Deviation(1593917502583_blob.png) 2.2.

For the sampling distribution, Mean is equal to the population mean.

Sample Standard Deviation of Sample(S)= population standard deviation(1593917502583_blob.png)/ N= 2.2/29= 0.40853

Standardizing the sampling distribution:

Z= S 71 T

Z stat for (x=137.8)= (137.8-136.9)/0.40853= 2.203022

Z stat for (x=138.1)=(138.1-136.9)/0.40853=2.937363

Using Z stat table,

P(2.203022<x<2.937363) = 0.012142

Л

Therefore , the probability that the average length of a randomly selected bundle of steel rods is between 137.8-cm and 138.1-cm.= 0.0121